Word Problem to solve - thank you
need 50 warheads and silos to complete project
Cost per warhead is 4 million dollars and silo cost is 6 million
What is the number of warheads and silos I can produce if the budget is 210 million
Need the following steps:
1. assign variables
2. write algebraic model
3. solve the model
4. answer
thanks
BB

Answer by tutorcecilia(2152)

x+y=50
Let x=warheads
Let y = silos
x+y=50
Than y=50-x
.
Each warhead (x) costs 4M; each silo (y) costs 6M; total cost is 210M
4x+6y=210 [substitues y=50-x]
4x+6(50-x)=210 [simplify and combine like-terms wherever possible]
4x+300-6x=210
-2x+300=210 [solve for the x=term]
-2x+300-300=210-300
-2x=-90
-2x/-2=-90/-2
x=45 warheads
.
So, x+y=50
45+y=50
y=5 silos
.
Check by using the original equation, plug-in the values:
4x+6y=210
4(45)+6(5)=210
180+30=210
210=210 [checks out]